# HARMONY: Hidden Activation Representations and Model Output-Aware Uncertainty Estimation for Vision-Language Models

### v1

@@ B.2 Training Strategy
We maintain two sets of data; calibration data and test data. Calibration data is further split into 80% and 20% split into training and validation data, respectively. For each model-dataset training, and every trainable scoring method, we perform hyper-parameter tuning of learning rate over { 5e-4, 5e-5, 5e-6}. We found 5e-5 lr to be working the best for most experiments. We use AUROC metric as our best model checkpoint selection critera for the methods. For the best model checkpoint, we report PRR, AURAC, Coverage at risk $10\%$ and $20\%$, and effective reliability. Further, for all the trainable methods, we use 20 epochs, and used early stopping; i.e, if validation auroc does not improve for 1K training steps, we stop the training. For MSF implementation, we follow the MLP architecture details from the official implementation of MSF (specifically, VisualBERT architecture experiments). We also keep other parameters such as optimizer (AdamW), learning rate scheduler (Warmup Cosine Scheduler) and batch size also the same. We use the same optimizer and learning rate for our method and LARS as well. Further, we use batch size of 32 for LARS and HARMONY. Since we are working on open-ended generations which can be of arbitrary length, therefore, we use zero padding to keep hidden representations of same length that is 128 for all the methods. For probability split in LARS, and our method, we use a bin count of 8.
B.3 Some Representative Examples
Figure 6: An illustration of selective prediction decisions on the A-OKVQA dataset with LLaVa-7b model. In the left-most example, the model generates an incorrect answer, yet both LARS [32] MSF [31] choose to answer based on their respective calibration thresholds. In the second example, LARS opts to answer, while MSF correctly abstains. In the right-most example, both methods abstain, whereas our approach makes the right prediction for each of these examples.

### v2

@@ A.2.3 Hyper-Parameters
We maintain two sets of data; calibration data and test data. Calibration data is further split into 80% and 20% split into training and validation data, respectively. For each model-dataset training, and every trainable scoring method, we perform hyper-parameter tuning of learning rate over { 5e-4, 5e-5, 5e-6}. We found 5e-5 lr works best for most experiments. We use the AUROC metric as our best model checkpoint selection criterion. We also show this in Figure 7 where we record AUROC and PRR on test set if the validation AUROC is >= the previous best. Finally, the best validation auroc step has been plotted by a yellow circle. For the best checkpoint, we report AUROC and PRR on the test set. We also use that checkpoint to report Coverage at risk $10\%$ and $20\%$, and effective reliability as reported in Table 2. Further, for all the trainable methods, we use 20 epochs, and used early stopping; i.e, if validation auroc does not improve for 1K training steps, we stop the training. For MSF implementation, we follow the MLP architecture details from the official implementation of MSF (specifically, VisualBERT architecture experiments). We also keep other parameters such as optimizer (AdamW), learning rate scheduler (Warmup Cosine Scheduler) and batch size also the same. We use the same optimizer and learning rate for our method and LARS as well. Further, we use batch size of 32 for LARS and HARMONY. Since we are working on open-ended generations which can be of arbitrary length, therefore, we use zero padding to keep hidden representations of same length that is 128 for all the methods. For probability split in LARS, and our method, we use a bin count of 8.
(a) AOKVQA - Validation AUROC
(b) AOKVQA - Test AUROC
(c) AOKVQA - Test PRR
(d) PathVQA - Validation AUROC
(e) PathVQA - Test AUROC
(f) PathVQA - Test PRR
Figure 7: Comparison of learnable scoring functions, MSF, LARS and HARMONY during training on the AOKVQA and PathVQA datasets using the Qwen-VL model. Validation AUROC is used to select the best checkpoint (highlighted with a yellow circle), and the corresponding Test AUROC and PRR scores are reported in the proceeding plots and Table 1.
A.2.4 Additional Results on Medical Domain
In Table 9, we also report the UE performance on the medical imaging domain dataset, PathVQA with Qwen-VL model. For this dataset and model as well, we find that the output probability based scoring functions perform significantly lower than the learnable functions as shown in 9. Further, HARMONY achieves state-of-the-art performance indicating its effective performance in medical domain.
Table 9: UE Performance on PathVQA
Qwen-VL
UE Method AUROC(%) PRR(%)
Length-Normalized Confidence 71.18 26.74
First Token Confidence 71.64 26.65
Self-Eval Confidence 71.91 30.57
Entropy 76.03 45.86
Semantic Entropy 63.72 15.42
Cluster Entropy 61.79 10.97
MSF 91.36 95.62
LARS 91.99 95.82
HARMONY [Ours] 92.92 96.30
(+0.93) (+0.68)
A.2.5 Computation
We perform all inference generations and evaluations on NVIDIA A100 40GB GPUs. For all inference generations and evaluations including black-box baselines, such as Semantic Entropy, we used two GPUs. Generation and labeling of the calibration data for one dataset containing  13k-16k samples and one model with LAVE labeling [manas2024improving] takes approximately 8-9 GPU-hours. Training a scoring function required at most two hours per run.